40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 27 04:36:46 2015, den, Update, LSC, als->pdh transition, prcl on 1f, alignment 
    Reply  Fri Mar 27 14:47:57 2015, Koji, Update, LSC, als->pdh transition, prcl on 1f, alignment CARM_transition.pdf
Message ID: 11179     Entry time: Fri Mar 27 14:47:57 2015     In reply to: 11177
Author: Koji 
Type: Update 
Category: LSC 
Subject: als->pdh transition, prcl on 1f, alignment 

Jenne and I interviewd Den this afternoon to make the things clear

- His "duty cycle" is not about the lengths of the lock stretch. He saids, the transition success probability is improved.

- For this improvement, the CARM transition procedure was modified to include turning on 1:20 (Z1P20) filter in CARM_A (i.e. ALS) once CARM_B (i.e. RF) dominates the loop in all frequency.

- I think this transition can be summarized like the attachment. At STEP4, the integration of the ALS is reduced. This actually does not change the stability of the servo as the servo stability is determined by the stability of the CARM_B loop. But this does further allow CARM_B to supress the noise. Or in other word, we can remove the noise coming from the CARM_A loop.

- The POP22 issue: Jenne has the trigger signal that is immune to this issue by adding some amount of POPDC for the trigger.
We can avoid the trigger issue by this technique. But if the issue is due to the true optical gain fluctuation, this may mean that the 11MHz optical gain is changing too much. This might be helped by PRC angular feedforward or RF 22MHz QPD at POP.

Attachment 1: CARM_transition.pdf  24 kB  Uploaded Fri Mar 27 17:01:12 2015  | Hide | Hide all
CARM_transition.pdf
ELOG V3.1.3-